OpenStudy (anonymous):

\[A~curve~is~given~by~the~equation~ y={2 \over 3}e^x+ {1 \over 3}e^{-2x}.\](a)Evaluate a definite integral to find the area between the curve, the x-axis and the lines x=0 and x=1, showing your working (b)Use calculus to determine whether the turning point at the point where x=0 is a maximum or minimum.

sam (.sam.):

\[\int\limits_{}^{}(\frac{2}{3}e ^{x}+\frac{1}{3}e ^{-2x})dx\]

sam (.sam.):

\[\huge [\frac{2}{3}e ^{x}+\frac{1}{3}\frac{1}{-2}e ^{-2x}]^{1}_{0}\] \[\huge [\frac{2}{3}e ^{x}+\frac{1}{-6}e ^{-2x}]^{1}_{0}\]

sam (.sam.):

For b), differentiate y \[\frac{dy}{dx}=\frac{2}{3}e ^{x}-\frac{2}{3}e ^{-2x}\]

sam (.sam.):

\[\frac{dy}{dx}=0\] \[\frac{2}{3}e ^{x}-\frac{2}{3}e ^{-2x}=0\] Solve for x

sam (.sam.):

To find whether is a max or min, \[\frac{d ^{2}y}{dx ^{2}}\] If value of \[\frac{d ^{2}y}{dx ^{2}}>0\] Then it is a minimum. If value of \[\frac{d ^{2}y}{dx ^{2}}<0\] Then it is a maximum.

OpenStudy (anonymous):

How do you differentiate it the second time?

sam (.sam.):

Use the differentiated equation differentiate again

OpenStudy (anonymous):

What will that give? \[{2 \over 3}e -{1\over12}e^{-2x}?\]

OpenStudy (anonymous):

*2/3 e^x* sorry

sam (.sam.):

\[\frac{d^{2}y}{dx ^{2}}=\frac{2}{3}e ^{x}+\frac{4}{3}e ^{-2x}\]

OpenStudy (anonymous):

Why?

sam (.sam.):

differentiate \[e ^{-2x}\] \[\frac{dy}{dx}=-2e ^{-2x}\]

sam (.sam.):